Circular No. 8639

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005kz AND 2005la Two supernova discoveries have been reported from unfiltered CCD

frames: 2005kz by T. Puckett and G. Sostero (via the 0.60-m automated supernova patrol telescope in Ellijay, GA; cf. IAUC 8635), and 2005la by R. Quimby and P. Mondol (via the 0.45-m ROTSE-IIIb telescope at the McDonald Observatory; cf. IAUC 8622).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005kz Dec. 1.03 19^{h}00^{m}49.^{s}82 +50^{o}53′01′′.8 18.7 15′′.6 E, 9′′.5 N
2005la Nov. 30.51 12 52 15.68 +27 31 52.5 17.6 6′′ W, 6′′ S

Additional approximate magnitudes for 2005kz in MCG +08-34-32: Nov. 1, [20.0; Dec. 1.99 UT, 18.2. Additional approximate magnitudes for 2005la in KUG 1249+278: 2004 Dec. 15, [19.0; 2005 Jan. 16, [19.0; Dec. 1.49, 17.6. Quimby adds that the position tabulated above for 2005la has uncertainty ± 0′′.8.

A. V. Filippenko and R. J. Foley, University of California, Berkeley; and T. Matheson, National Optical Astronomy Observatory, report that inspection of a CCD spectrogram (range 330–930 nm), obtained on Dec. 4 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2005kz is of type Ic, perhaps 1 week after maximum brightness, with a peculiar spectrum resembling those of SN 1998bw (Patat et al. 2001, Ap.J. 555,

900) and SN 2002ap (Foley et al. 2003, PASP 115, 1220). The Ca II near-infrared triplet has an unusually broad absorption trough, thoroughly blended with the O I 777.4-nm line.

Filippenko, Foley, and Matheson add that inspection of a CCD spectrogram (range 400–900 nm), obtained by J. Cooke and J. C. Berrier (University of California, Irvine) on Dec. 3 UT with the Keck II 10-m telescope (+ DEIMOS), shows that SN 2005la is a peculiar type-II/Ib supernova: “In addition to relatively weak, narrow hydrogen Balmer lines having P-Cyg profiles, the spectrum exhibits narrow He I lines with P-Cyg profiles dominated by the emission component. The recession velocity of the host galaxy is 5570 km/s, based on narrow Hα from nearby H II regions. The supernova expansion velocity, determined from the minima in the absorption components, is ∼ 1700 km/s. The FWHM of the emission components is ∼ 1800 km/s. We have previously not encountered a supernova spectrum having these characteristics. Further monitoring of the object is encouraged.”

2005 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green